Consider a Boeing 747 flying at a standard altitude of 36,000 ft. The pressure at a point on the wing is 400 lb/ft2. Assuming isentropic flow over the wing, calculate the temperature at this point


1
Expert's answer
2021-03-12T05:42:47-0500


We know that isentropic process means constant entropy process so, it may also be adiabatic process for reversible condition without heat transfer.

So for adiabatic process we know the relation between temeprature and pressure as

In standard condition pressure is "P_1= 476 \frac{lb}{ft^2}" and temperature is 391oR, and pressure above altitude is 400 lb/ft2,

Now,

"T_2= T_1(\frac{P_1}{P_2})^{\frac{\gamma -1}{\gamma}}"


"T_2= 391(\frac{400}{476})^{\frac{1.4 -1}{1.4}}"


T2= 372. 04 oR

This is temperature as altitude of 36,000 ft